# Cycles in spherical Deligne complexes and application to $K(\pi,1)$-conjecture for Artin groups

### v1

@@ 1.3 Comment on the proofs
Let $A_{S}$ be an Artin group with generating set $S$. Let $S^{\prime}\subset S$. The $(S,S^{\prime})$-relative Artin complex $\Delta_{S,S^{\prime}}$ is defined to be the induced subcomplex of the Artin complex $\Delta_{S}$ of $A_{S}$ spanned by vertices of type $\hat{s}$ with $s\in S^{\prime}$.
An Artin group $A_{S}$ is almost spherical if its presentation graph is complete, it is not spherical, but for each $s\in S$, $A_{\hat{s}}$ is spherical. A subset $T\subset S$ is almost spherical if $A_{T}$ is almost spherical. Almost spherical Artin groups are classified: they are either affine, or  one of the hyperbolic Lannér types. If $A_{S}$ is almost spherical, then the link of each vertex is isomorphic to the Artin complex of a spherical Artin group, which is not contractible by previous discussion, thus the deformation retraction procedure does not work for $\Delta_{S}$. Thus, these Artin complexes and their relative versions serve as our “cores”.  We conjecture that the cores are contractible. The following conjecture actually implies the $K(\pi,1)$-conjecture for all Artin groups [Hua23].
Conjecture 1.13 ([Hua23]).
Suppose $S^{\prime}\subset S$, and $S^{\prime}$ is almost spherical. Suppose the Dynkin diagrams for $S^{\prime}$ and $S$ are connected and do not have $\infty$-labeled edges. Then $\Delta_{S,S^{\prime}}$ is contractible.
We plan to metrize $\Delta_{S,S^{\prime}}$ as a “non-positive curved metric space” to show contractibility.  There is a natural piecewise Euclidean or piecewise hyperbolic metric we can put on $\Delta_{S,S^{\prime}}$ [Hua23] (see Section A), and the resulting metric space is conjecturally CAT$(0)$ or CAT$(-1)$, which gives contractibility. However, the feasibility of proving this conjecture is not clear due to the difficulty of verifying CAT$(0)$ for higher dimensional complexes.
A different metric on $\Delta_{S,S^{\prime}}$, discussed in [Hua23], is conjectured to make $\Delta_{S,S^{\prime}}$ a metric space with convex geodesic bicombing in the sense of [DL15], hence contractible. We have a candidate for such metric if $S^{\prime}$ does not belong to the following examples: $\{\widetilde{F}_{4},\widetilde{E}_{6},\widetilde{E}_{7},\widetilde{E}_{8},[3,5,3],[5,3,3,3]\}$. See Section A for details of the proposed metric in each case.
Combining with the work of Haettel [Hae22], in order to verify the alternative metric in Section A has the desired property which leads to the contractibility of the complex, it suffices to understand cycles of length $\leq 6$ in the link of each vertex of $\Delta_{S,S^{\prime}}$ when $\dim(\Delta_{S,S^{\prime}})\geq 3$, and cycles of length $\leq 12$ in the vertex links of $\Delta_{S,S^{\prime}}$ when $\dim(\Delta_{S,S^{\prime}})=2$.
Thus, the core of this strategy is to understand cycles up to a certain length in relative Artin complexes. This turns out to be the most challenging part, as understanding an $n$-cycle in $\Delta_{S,S^{\prime}}$ is equivalent to understanding the solution set of the equation
$w_{1}w_{2}\cdots w_{n}=1$ (1.14)
in $A_{S}$, subject to the constraint that each $w_{i}$ is contained in a suitable parabolic subgroup of $A_{S}$. Thus the whole plan relies on the feasibility of solving (1.14), which is not clear at all - even the word problem for most of these Artin groups is open.
Our plan is to use a propagation argument to reduce the study of cycles in general relative Artin complexes to Artin complexes associated with spherical Artin groups. For a vertex $v$ of type $\hat{s}$ in $\Delta_{S,S^{\prime}}$, $\operatorname{lk}(v,\Delta_{S,S^{\prime}})\cong\Delta_{T,T^{\prime}}$ with $T=S\setminus\{s\}$ and $T^{\prime}=S^{\prime}\setminus\{s\}$. If $A_{T}$ is already spherical, then the reduction is done. If $A_{T}$ is not spherical, then we can find $T^{\prime\prime}\subset T$ such that $A_{T^{\prime\prime}}$ is almost spherical. If we understand, say, 6-cycles in the vertex links of $\Delta_{T,T^{\prime\prime}}$, then we can conclude that $\Delta_{T,T^{\prime\prime}}$ is non-positively curved in an appropriate sense. While $T^{\prime\prime}$ might not equal $T^{\prime}$, $A_{T}$ acts on both $A_{T,T^{\prime\prime}}$ and $A_{T,T^{\prime}}$, this allows us to encode a 6-cycle in $\Delta_{T,T^{\prime}}$ as six nice subspaces of $\Delta_{T,T^{\prime\prime}}$, and use the non-positive curvature of $\Delta_{T,T^{\prime\prime}}$ to analyze the configuration of these subspaces. This gives potentially a way to use our knowledge of 6-cycles in the vertex links of $\Delta_{T,T^{\prime\prime}}$ associated to a smaller Artin group $A_{T}$ to understand 6-cycles in the vertex links of $\Delta_{S,S^{\prime}}$ associated with a larger Artin group $A_{S}$. We can keep doing this until $A_{T}$ is spherical.  [^1] An important, though a bit more technical point - the properties of cycles that propagate properly under this procedure are different from the existing links in the literature (usually one has to require more), however, we expect that it still concerns cycles of the same length.
Now we are reduced to studying cycles of length $\leq 6$ in $\Delta_{S}$ with $S$ spherical, and cycles of length $\leq 12$ in $\Delta_{S}$ with $S$ spherical and $\Delta_{S}$ being 1-dimensional. The later is already understood. The former is still highly non-trivial.

### v3

@@ 2.4 Two contractibility criterion for simplicial complexes
The interest in the bowtie-free condition lies in the following observation.
Lemma 2.12.
[BM10, Proposition 1.5] If $P$ is a bowtie free graded poset, then any pair of elements in $P$ with a lower bound have a join, and any pair of elements in $P$ with an upper bound have a meet.
Let $P$ be a bounded graded poset. Then $P$ is lattice if it is bowtie free.
Definition 2.13.
A poset $P$ is upward flag if any three pairwise upper bounded elements have an upper bound. A poset is downward flag if any three pairwise lower bounded elements have a lower bound. A poset is flag if it is both upward flag and downward flag.
Definition 2.14.
  1. 1.
    $x,y,z$ are not maximal element in $P$;
  2. 2.
    each pair in $\{x,y,z\}$ have a upper bound in $P$ which is not maximal.
Let $S=\{s_{1},s_{2},\ldots,s_{n}\}$. A simplicial complex $X$ is of type $S$ if all the maximal simplices of $X$ has dimension $n-1$ and there is a type function $\operatorname{Type}$ from the vertex set of $X$ to $\{\hat{s}_{1},\hat{s}_{2},\ldots,\hat{s}_{n}\}$ such that $\operatorname{Type}(x)\neq\operatorname{Type}(y)$ whenever $x$ and $y$ are adjacent vertices of $X$. This labeling induces a bijection between $\{\hat{s}_{1},\hat{s}_{2},\ldots,\hat{s}_{n}\}$ and the vertex set of each maximal simplex of $X$.
Note that if $A_{S^{\prime}}$ is an Artin group, and $S$ is a subset of the set of generators $S^{\prime}$, then the relative Artin complex $\Delta_{S^{\prime},S}$ is a simplicial complex of type $S$. We will be interested in more general simplicial complexes of type $S$, for some $S$ not necessarily made of generators of some Artin groups.
Definition 2.15.
Now we discuss a situation when the relation $<$ on $V$ is a poset.
Definition 2.16.
An induced subgraph $\Lambda^{\prime}$ of $\Lambda$ is admissible if for any node $x\in\Lambda^{\prime}$, if $x_{1},x_{2}\in\Lambda^{\prime}$ are in different connected components of $\Lambda^{\prime}\setminus\{x\}$, then they are in different components of $\Lambda\setminus\{x\}$.
Lemma 2.17.
([Hua24a, Lemma 6.6]) Suppose $\Lambda^{\prime}$ is an admissible linear subgraph of the Coxeter diagram $\Lambda$ of $A_{S}$ and suppose the consecutive nodes of $\Lambda^{\prime}$ are $S^{\prime}=\{s_{i}\}_{i=1}^{n}$. Let $\Delta$ be the $(\Lambda,\Lambda^{\prime})$-relative Artin complex. Let $V$ be the vertex set of $\Delta$. We order $S^{\prime}$ such that it is compatible with one of the two linear orders on $\Lambda^{\prime}$. Then the induced relation $<$ on $V$ is a graded poset.
Now we are ready to state Haettel’s result. Let $S=\{s_{1},\ldots,s_{n}\}$ with a total order $s_{1}<s_{2}<\cdots<s_{n}$. The vertex set of a simplicial complex of type $S$ is endowed with the relation induced from such total order on $S$. The following is a consequence of [Hae22, Section 4.3, Theorem B] and [Hae21, Theorem 1.15].
Theorem 2.18.
[Hae21, Hae22] Let $X$ be a simplicial complex of type $S$. Assume that
  1. 1.
    $X$ is simply connected;
  2. 2.
    the relation $<$ on the vertex set $V$ of $X$ is a partial order;
  3. 3.
  4. 4.
Then $X$ is contractible, and $X$ admits a piecewise $\ell^{\infty}$ injective metric.
The graph $Y$ in the above theorem is called the thickening of $X$. Given a simplicial graph $Z$ endowed with the path metric such that each edge has length $1$, a combinatorial ball in $Z$ is the collection of vertices in the metric ball $B(x,r)$ of radius $r$ centered at a vertex $x$. The graph $Z$ is Helly if whenever a collection of combinatorial balls in $Z$ have non-empty pairwise intersection, then the common intersection of these balls is non-empty. Helly graphs are closely related to the notion of injective metric spaces, which are continuous versions of Helly graphs. A geodesic metric space is injective if whenever a collection of closed metric balls in the space have non-empty pairwise intersection, then the common intersection of these balls is non-empty.
Now we discuss a variation of Theorem 2.18. Put a cyclic order $s_{1}<s_{2}<\cdots s_{n}<s_{1}$ on $S$. For each vertex $x$ in $X$ of type $\hat{s}_{i}$, we consider a relation $<_{x}$ in $\operatorname{lk}(x,X)$ as follows. We identify vertices in $\operatorname{lk}(x,X)$ as vertices in $X$ which are adjacent to $x$. For each $s_{i}\in S$, this cyclic order induces an order on $S\setminus\{s_{i}\}$ by declaring $s_{i+1}<\cdots<s_{n}<s_{1}<\cdots s_{i-1}$. For $y,z\in\operatorname{lk}(x,X)$, define $y<_{x}z$ if $y$ is adjacent to $z$ and $\operatorname{Type}(y)<\operatorname{Type}(z)$ in $S\setminus\{s_{i}\}$. The following is a consequence of [Hae22, Section 4.2, Theorem A].
Theorem 2.19.
  1. 1.
    $X$ is simply-connected.
  2. 2.
  3. 3.
2.5 Subdivision of some relative Artin complexes
In this subsection, we explain the connection between Conjecture 2.6 and the two contractibility criteria in Section 2.4. This is done by adjusting the procedure of handling Euclidean buildings and certain Artin complexes in Haettel’s work [Hae21] to certain relative Artin complexes.
Definition 2.20.
Suppose $\Lambda^{\prime}$ is an admissible linear subgraph of the Coxeter diagram $\Lambda$ of $A_{S}$ with consecutive nodes of $\Lambda^{\prime}$ being $\{s_{i}\}_{i=1}^{n}$. We say the $(\Lambda,\Lambda^{\prime})$-relative Artin complex $\Delta_{\Lambda,\Lambda^{\prime}}$ is bowtie free if the poset in Lemma 2.17 is bowtie free. We $\Delta_{\Lambda,\Lambda^{\prime}}$ is flag or weakly flag, if the poset in Lemma 2.17 is flag or weakly flag respectively. Note that these definitions do not depend on the choice of the linear order on $\Lambda^{\prime}$.
Definition 2.21.
Suppose $A_{\Lambda}$ is an Artin group whose Coxeter diagram $\Lambda$ contain induced subgraph $\Lambda^{\prime}$ such that $\Lambda^{\prime}$ is a copy of Coxeter diagram of type $\widetilde{B}_{n}$, though possibly with different edge labeling. Let $\{b_{i}\}_{i=1}^{n}$ be nodes of $\Lambda^{\prime}$ as in Figure 3 left.
Let $\Delta=\Delta_{\Lambda,\Lambda^{\prime}}$ be the associated relative Artin complex. We subdivide each edge of $\Delta$ connecting a vertex of type $\hat{b}_{1}$ and a vertex of type $\hat{b}_{2}$. We say the middle point of such edge is of type $m$. Cut each top dimensional simplex in $\Delta$ into two simplices along the codimensional 1 simplex spanned by vertices of type $m$ and $\{\hat{b}_{i}\}_{i=3}^{n+1}$. This gives a new simplicial complex, which we denoted by $\Delta^{\prime}$. Define a map $t$ from the vertex set $V\Delta^{\prime}$ of $\Delta^{\prime}$ to $\{1,2,\ldots,n,n+1\}$ by sending vertices of type $\hat{b}_{1},\hat{b}_{2}$ to $1$, vertices of type $m$ to $2$, vertices of type $\hat{b}_{i}$ to $i$ for $i\geq 3$. We will then view $\Delta^{\prime}$ as a simplicial complex of type $S=\{1,2,\ldots,n+1\}$. We define a relation $<$ on $V\Delta^{\prime}$ as follows. For $x,y\in V\Delta^{\prime}$, $x<y$ if $x$ and $y$ are adjacent and $t(x)<t(y)$. The simplicial complex $\Delta^{\prime}$, together with the relation $<$ on its vertex set, is called the $(b_{1},b_{2})$-subdivision of $\Delta_{\Lambda,\Lambda^{\prime}}$.
Figure 3: Diagrams of type $\widetilde{B}$ and $\widetilde{D}$.
Lemma 2.22.
Proof.
Lemma 2.23.
([Hua24a, Lemma 6.2]) Suppose $|S^{\prime}|\geq 3$. Then the relative Artin complex $\Delta_{S,S^{\prime}}$ is simply-connected.
Proposition 2.24.
Suppose $A_{\Lambda}$ is an Artin group whose Coxeter diagram $\Lambda$ contain induced subgraph $\Lambda^{\prime}$ such that $\Lambda^{\prime}$ is a copy with Coxeter diagram of type $\widetilde{B}_{n}$, though possibly with different edge labeling. Let $\{b_{i}\}_{i=1}^{n+1}$ be nodes of $\Lambda^{\prime}$ as in Figure 3 left. Suppose $n\geq 3$. Suppose $\Lambda^{\prime}$ is an admissible subgraph of $\Lambda$.
  1. 1.
  2. 2.
Proof.
Definition 2.25.
Suppose $A_{\Lambda}$ is an Artin group whose Coxeter diagram $\Lambda$ contain induced subgraph $\Lambda^{\prime}$ such that $\Lambda^{\prime}$ is a copy with Coxeter diagram of type $\widetilde{D}_{m}$ with $m\geq 4$, though possibly with different edge labeling. Let $a_{1},a_{2},\{b_{i}\}_{i=1}^{n},c_{1},c_{2}$ be nodes of $\Lambda^{\prime}$ as in Figure 3 right.
Let $\Delta=\Delta_{\Lambda,\Lambda^{\prime}}$ be the associated relative Artin complex. We subdivide each edge of $\Delta$ connecting a vertex of type $\hat{a}_{1}$ (resp. $\hat{c}_{1}$) and a vertex of type $\hat{a}_{2}$ (resp. $\hat{c}_{2}$). We say the middle point of such edge is of type $\hat{a}$ (resp. $\hat{c}$). Cut each top dimensional simplex in $\Delta$ into four simplices whose vertex set is of type $\{\hat{a}_{i},\hat{a},\hat{b}_{1},\ldots,\hat{b}_{n},\hat{c},\hat{c}_{j}\}_{1\leq i,j\leq 2}$. This gives a new simplicial complex, which we denoted by $\Delta^{\prime}$. Define a map $t$ from the vertex set $V\Delta^{\prime}$ of $\Delta^{\prime}$ to $\{1,\ldots,n+4\}$ by sending vertices of type $\hat{a}_{i},\hat{a},\hat{b}_{1},\ldots,\hat{b}_{n},\hat{c},\hat{c}_{j}$ to $41,2,3,\ldots,n+2,n+3,n+$ 4 respectively. We will then view $\Delta^{\prime}$ as a simplicial complex of type $S=\{1,\ldots,n+4\}$. We define a relation $<$ on $V\Delta^{\prime}$ as follows. For $x,y\in V\Delta^{\prime}$, $x<y$ if $x$ and $y$ are adjacent and $t(x)<t(y)$. Similar to Lemma 2.22, we know $(V\Delta^{\prime},\leq)$ is a poset, under the additional assumption that $\Lambda^{\prime}$ is admissible in $\Lambda$.
Proposition 2.26.
Suppose $A_{\Lambda}$ is an Artin group whose Coxeter diagram $\Lambda$ contains induced subgraph $\Lambda^{\prime}$ such that $\Lambda^{\prime}$ is a copy with Coxeter diagram of type $\widetilde{D}_{m}$ with $m\geq 4$, though possibly with different edge labeling. Let $a_{1},a_{2},\{b_{i}\}_{i=1}^{n},c_{1},c_{2}$ be nodes of $\Lambda^{\prime}$ as in Figure 3 right. Suppose $\Lambda^{\prime}$ is an admissible subgraph of $\Lambda$.
  1. 1.
  2. 2.
Corollary 2.27.
Suppose $\Lambda$ is a connected Coxeter diagram, and $\Lambda^{\prime}$ is an admissible subgraph of $\Lambda$ in the sense of Definition 2.16. Suppose $\Lambda^{\prime}$ is almost spherical with vertex set $S^{\prime}$ and $\Lambda$ has vertex set $S$. Suppose $\Lambda^{\prime}$ is not of type $\{\widetilde{F}_{4},\widetilde{E}_{6},\widetilde{E}_{7},\widetilde{E}_{8},[3,5,3],[5,3,3,3]\}$. Then there is a criterion only involving cycles in the 1-skeleton in the link of each vertex of $\Delta_{S,S^{\prime}}$ such that as long as such link criterion is satisfied, then Conjecture 2.6 holds true, in particular $\Delta_{S,S^{\prime}}$ is contractible. More precisely,
  1. 1.
    if $S^{\prime}$ dominates type $\widetilde{A}_{n}$ (in which case the Coxeter diagram of $S^{\prime}$ must be a cycle, hence elements in $S^{\prime}$ has a natural cyclic order), then we need to check $\Delta_{S,S^{\prime}}$, viewed as a simplicial complex of type $S^{\prime}$, satisfies the assumption 3 of Theorem 2.19;
  2. 2.
    if $S^{\prime}$ dominates type $\widetilde{C}_{n}$ (in which case the Coxeter diagram of $S^{\prime}$ is a line, hence $S^{\prime}$ has two linear orders), then we need to check if $S^{\prime}$ is equipped with one of these linear orders, then $\Delta_{S,S^{\prime}}$ satisfies the assumptions 3 and 4 of Theorem 2.18;
  3. 3.
    if $S^{\prime}$ dominates type $\widetilde{B}_{n}$ (in which case the Coxeter diagram of $S^{\prime}$ is isomorphic to Figure 3 left), then we need to check assumptions of Proposition 2.24 are satisfied;
  4. 4.
    if $S^{\prime}$ dominates type $\widetilde{D}_{n}$ (in which case the Coxeter diagram of $S^{\prime}$ is isomorphic to Figure 3 right), then we need to check assumptions of Proposition 2.26 are satisfied;
  5. 5.
    if $S^{\prime}=\{a,b,c\}$ dominates type $\widetilde{G}_{2}$ with $m_{ab}\geq 3$ and $m_{bc}\geq 6$, then we need to check $\operatorname{lk}(x,\Delta_{S,S^{\prime}})$ is a graph with girth $\geq 6$ whenever $x$ is of type $\hat{c}$ and $\operatorname{lk}(x,\Delta_{S,S^{\prime}})$ is a graph with girth $\geq 12$ whenever $x$ is of type $\hat{a}$.
Proof.
Assertions 1 and 2 follow from Lemma 2.23, Lemma 2.22, Theorem 2.19 and Theorem 2.18. Assertions 3 and 4 follows from Proposition 2.24 and Proposition 2.26. For Assertion 5, we metrize triangles with flat metric, with angle $\pi/2$ at vertices of type $\hat{b}$, angle $\pi/6$ at vertices of type $\hat{a}$ and angle $\pi/3$ at vertices of type $\hat{c}$. The assumptions imply $\Delta_{S,S^{\prime}}$ is locally CAT$(0)$, hence we are done by Lemma 2.23. ∎